I. A family buys a house worth $326,000. They pay $75,000 deposit and take a mortgage for the balance at J12=9% p.a. to be amortized over 30 years with monthly payments.

A. Find the value of the mortgage on their house?
B. Find the value of the monthly payment?
C. Find the loan outstanding after making 20 payments?
D. Find the principal repaid in the 21st payment?
1
Expert's answer
2020-06-04T19:05:34-0400

A. value of the mortgage = 326,000 - 75,000 = 251,000

B. monthly payment = 251,000*(0.09/12)/((1+0.09/12)^(12*30)-1)=137.1027

C. Principal Outstanding after making 20 payments = 137.1027*(1-(1+0.09/12)^(-20))/(0.09/12)= 2537.49947166

D. Principal paid in 21st Payment = 251,000-2537.49947166 = 248462.500528